Investigating - Ticket#400113542747. Torso coil failed QA/multiple coil element issues out. Will coordinate with Siemens on a good time to assess. Scanner is operational
Nov 18, 2025 - 14:01 EST
Identified - Siemens believes this is a protocol issue. They suggest rebuilding the sequence and if the problem still persists after that, then they would have to contact someone in Germany to take a look at it.
Nov 14, 2025 - 15:41 EST
Update - Tech support investigating the issue and will update 11/14.
Nov 13, 2025 - 17:48 EST
Investigating - Ticket# 400113534666. Repeatedly experiencing 3D shim errors while shimming for VISTA sequence. Consulting Siemens tech support for help to investigate. Scanner is operational
Nov 13, 2025 - 11:01 EST
Monitoring - Laser light worked this am, but not this afternoon. Previous ticket#400113533682 is still open. Siemens does not have an Eta as of 11/14/25 @ 2:35pm
Nov 14, 2025 - 14:46 EST
Investigating - There is a RF noise spike in the higher bandwidth range when running CAMRIS QA on the 20-channel head coil. An image artifact was seen on a 3D Flash sequence at higher bandwidths. We are currently investigating.
Apr 07, 2025 - 15:29 EDT
Identified - We are awaiting a replacement monitor for SC3T, and in the interim the monitor from SC7T has been moved to SC3T. We expect a replacement monitor will be available this summer, and will provide a more specific update once we have additional information from the manufacturer.
Apr 24, 2025 - 14:40 EDT
Investigating - We are currently investigating this issue.
Apr 24, 2025 - 14:39 EDT

About This Site

This is the status page for CAMRIS-maintained MRI scanners and ancillary equipment.

To report issues, please post in the #technical-issues channel of the UPenn MRI Slack (http://upennmri.slack.com). Reports will be triaged, and this page updated as necessary.

Stellar Chance 3T Operational
90 days ago
99.53 % uptime
Today
MRI Scanner ? Operational
90 days ago
92.55 % uptime
Today
Visual Stimulus Presentation ? Operational
90 days ago
100.0 % uptime
Today
Audio Stimulus Presentation Operational
90 days ago
100.0 % uptime
Today
Response Device System ? Operational
90 days ago
100.0 % uptime
Today
32-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
64-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
20-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
15-Channel Knee Coil Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring ? Operational
90 days ago
100.0 % uptime
Today
Eye Tracker Operational
90 days ago
100.0 % uptime
Today
Rico - Linux Host Operational
90 days ago
100.0 % uptime
Today
In-bore Camera Operational
90 days ago
100.0 % uptime
Today
FIRMM System Operational
90 days ago
100.0 % uptime
Today
Optical trigger Operational
90 days ago
100.0 % uptime
Today
RTexport Operational
90 days ago
100.0 % uptime
Today
Spine Coil Operational
90 days ago
100.0 % uptime
Today
HUP6 Operational
90 days ago
99.92 % uptime
Today
MRI Scanner ? Operational
90 days ago
99.12 % uptime
Today
Visual Stimulus Presentation ? Operational
90 days ago
100.0 % uptime
Today
Audio Stimulus Presentation Operational
90 days ago
100.0 % uptime
Today
Response Device System ? Operational
90 days ago
100.0 % uptime
Today
32-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
64-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring ? Operational
90 days ago
100.0 % uptime
Today
Eye Tracker Operational
90 days ago
100.0 % uptime
Today
Napoleon - Linux Host Operational
90 days ago
100.0 % uptime
Today
Optical trigger Operational
90 days ago
100.0 % uptime
Today
RTexport Operational
90 days ago
100.0 % uptime
Today
Stellar Chance 7T Major Outage
90 days ago
92.86 % uptime
Today
MRI Scanner ? Operational
90 days ago
99.51 % uptime
Today
pTx System ? Operational
90 days ago
100.0 % uptime
Today
Visual Presentation System ? Major Outage
90 days ago
0.52 % uptime
Today
Audio Stimulus Presentation Operational
90 days ago
100.0 % uptime
Today
Response Device System ? Operational
90 days ago
100.0 % uptime
Today
1Tx32Rx Head Coil Operational
90 days ago
100.0 % uptime
Today
8Tx32Rx Head Coil Operational
90 days ago
100.0 % uptime
Today
8-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
28-Channel Knee Coil Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring ? Operational
90 days ago
100.0 % uptime
Today
Eye Tracker Operational
90 days ago
100.0 % uptime
Today
Pedro - Linux Host Operational
90 days ago
100.0 % uptime
Today
In-bore Camera Operational
90 days ago
100.0 % uptime
Today
RTexport Operational
90 days ago
100.0 % uptime
Today
PAV10 Operational
90 days ago
99.96 % uptime
Today
MRI Scanner ? Operational
90 days ago
99.74 % uptime
Today
64-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring Operational
90 days ago
100.0 % uptime
Today
Response Device System ? Operational
90 days ago
100.0 % uptime
Today
Audio Stimulus Presentation ? Operational
90 days ago
100.0 % uptime
Today
32-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
20-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
HUP5 Operational
90 days ago
100.0 % uptime
Today
MRI Scanner ? Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring Operational
90 days ago
100.0 % uptime
Today
MRD1 Operational
90 days ago
100.0 % uptime
Today
MRI Scanner ? Operational
90 days ago
100.0 % uptime
Today
20-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
64-Channel Head Coil Operational
90 days ago
100.0 % uptime
Today
MRD2 Operational
90 days ago
100.0 % uptime
Today
MRI Scanner ? Operational
90 days ago
100.0 % uptime
Today
Physiological Monitoring Operational
90 days ago
100.0 % uptime
Today
MRD3 Operational
90 days ago
100.0 % uptime
Today
MRI Scanner ? Operational
90 days ago
100.0 % uptime
Today
PBR Calendar Operational
90 days ago
100.0 % uptime
Today
Flywheel Reaper ? Operational
90 days ago
100.0 % uptime
Today
CAMRIS Service Desk ? Operational
90 days ago
100.0 % uptime
Today
Teamviewer Remote Access Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Nov 19, 2025

No incidents reported today.

Nov 18, 2025
Resolved - Engineer was able to place a new mars unit. The cmrr bold sequences were tested and ran fine. System is operational.
Nov 18, 15:25 EST
Update - Engineer to arrive on site ~3pm. Continuing to have measurement recon errors
Nov 18, 14:05 EST
Investigating - #400113542505. Tech support dispatching Siemens engineer.
Nov 18, 12:13 EST
Nov 17, 2025

No incidents reported.

Nov 16, 2025

No incidents reported.

Nov 15, 2025

No incidents reported.

Nov 14, 2025
Resolved - Per Siemens note, the software has been reloaded on MARS and system is up.
Nov 14, 17:17 EST
Investigating - Tech support to dispatch an engineer out tomorrow (11/14) AM. Issue with the MARS again. Ticket# 400113536207. System is not operational
Nov 13, 20:29 EST
Resolved - Laser light issue is rectified.
Nov 14, 08:44 EST
Update - Engineer ETA 11/13 around 6pm to fix the issue.
Nov 13, 17:49 EST
Investigating - PAV10 laser light is dim. Called into Siemens for an engineer to replace ticket#400113533682. System is still operational, but users may need to turn the room lights off to center appropriately. Will update ETA for engineer.
Nov 12, 18:26 EST
Nov 13, 2025
Nov 12, 2025
Resolved - Tristate completed the work on the chiller piping.
Nov 12, 08:13 EST
Monitoring - Tristate filled the refrigerant in circuit 1, adjusted the water regulating valve, and checked the charge on circuit 2. System was able to regulate temperature after several phantom scans. System is now functional. We will resume regular operations at HUP6 tomorrow.
Sep 30, 14:16 EDT
Update - Tristate reassessed the chiller system: circuit 1 refrigerant leak - causing the compressors to shut off and the machine to overheat. Tristate advised to not scan. A technician will be out tomorrow to replace the refrigerant and fix the system.
Sep 29, 19:24 EDT
Investigating - Scanner keeps showing stimulation as being almost exceeded after a half hour of scanning. Last week, Siemens, maintenance and Tristate came to look at this issue. New ticket placed through Siemens. Ticket 400113457592.
Sep 29, 10:24 EDT
Nov 11, 2025

No incidents reported.

Nov 10, 2025

No incidents reported.

Nov 9, 2025

No incidents reported.

Nov 8, 2025

No incidents reported.

Nov 7, 2025

No incidents reported.

Nov 6, 2025

No incidents reported.

Nov 5, 2025
Resolved - Internet connectivity has been resolved.
Nov 5, 09:33 EST
Update - IT has fixed todays issue and continues to monitor the network.
Aug 29, 14:23 EDT
Update - We are continuing to monitor for any further issues.
Aug 29, 14:21 EDT
Update - No Network Connection on HUP 6 again today. Manually typing in patient info because the worklist won't connect with PennChart. IT ticket has been open for two weeks as this is an ongoing issue. Ticket Number IR9887454. Comment added this morning for IT.
Aug 29, 09:28 EDT
Update - We are continuing to monitor for any further issues.
Aug 29, 09:26 EDT
Monitoring - We have changed ports for the network cable which has changed the speed of the network from 100Mbs to 1Gb. This seems to have fixed the network speed for sending images to Sectra and Flywheel. As of right now, we have caught up on sending all images to flywheel and sectra. Please make sure all of your images are being seen in flywheel within the last two weeks. We will still be monitoring this situation for the next couple of days to make sure this new port is reliable.
Aug 28, 09:44 EDT
Update - The connectivity issues to Sectra and Flywheel are still on going. We are slowly sending back dated studies to flywheel. As of right now, everything up to 8/19 should have made it to flywheel. If you do not see your studies, please reach out.
Aug 27, 08:58 EDT
Update - Due to the ongoing connectivity issues on HUP6, we recommend that study teams collect image data directly onto a hard drive or flash drive at the scanner. Please utilize the ticket system: https://camris.atlassian.net/servicedesk/customer/portals to schedule technologist assistance or call HUP6 215-662-3057.

Please note, most scan acquired from Monday, 9/18 to present have not successfully transferred to Flywheel.
UPHS IT and Siemens are actively working together to resolve the issue. CAMRIS will continue to provide updates as progress is made. Please check Statuspage and Slack to stay up to date!

Aug 25, 15:01 EDT
Investigating - Per IT recommendation, NIC replacement requested through Siemens. Ticket number 400113400035.
Aug 25, 10:48 EDT
Resolved - The spiking seen was protocol related.
Nov 5, 09:32 EST
Monitoring - Cold head compressor was topped off with gas and the knocking sound went away. The gradient spike artifact was only reproducible with the 32CH coil so the engineer cleaned the plugs and the spiking went away. He has ordered new plugs and will install when they arrive.
Oct 2, 13:14 EDT
Update - We are continuing to investigate this issue.
Oct 2, 13:11 EDT
Update - Siemens engineer scheduled Thursday 3pm - will also evaluate the cold head.
Oct 1, 12:06 EDT
Investigating - Ticket # 400113460818. In process of scheduling a time for Siemens engineer to evaluate. (aiming for Thursday, 10/2 afternoon) System is still operational
Sep 30, 14:44 EDT